When sensors in a car detect a collision, they cause the reaction of sodium azide, NaN3, which generates nitrogen gas to fill the air bags within 0.03 s. 2NaN3(s)→2Na(s)+3N2(g) How many liters of N2 are produced at STP if the air bag contains 101 g of NaN3? Express your answer with the appropriate units.
Answer #1

molar mass of NaN3 = 65.0099 gm/mole

then 101 gm of NaN3 = 101 / 65.0099 = 1.55 mole

According to reaction 2 mole of NaN3 produce 3 mole of N2 then 1.55 mole of NaN3 produce 1.55 X 3 / 2 = 2.325 mole of N2

at STP 1 mole gas occupy volume = 22.414 liter then 2.325 mole of N2 gas occupu volume = 2.325 X 22.414 = 52.1 liter

52.1 Liter of N2 produced.